OpenStudy (anonymous):

Solve the system by graphing.If the system is inconsistent and has no solution,state this .If the system is dependent,write the form of the solution for any real number x. 4x+5y=10 y=-3/5x+3

OpenStudy (e.mccormick):

OK, so what did you get when you graphed them, or do you need help with that?

OpenStudy (anonymous):

I just solved it and this is what I got . (x,-2/3+2)

OpenStudy (e.mccormick):

Ummm... there can't be an x in the solution.

OpenStudy (anonymous):

Can you help me out if I am right or wrong and also I do not know how to graph 2 lines so help me out

OpenStudy (e.mccormick):

I find that for many graphing problems, an easy way to get a rough graph is to find the intercepts. So I take \(4x+5y=10\) and I put in 0 for x to find the y intercept. \(4(0)+5y=10\implies 5y=10 \implies y=2\) That means the point (0,2) is on the graph. Then I take \(4x+5y=10\) and I put in 0 for y to find the x intercept. \(4x+5(0)=10\implies 4x=10 \implies x=\frac{5}{2}\) So \((\frac{5}{2},0)\) is also on that line. Then I draw a line through that line.

OpenStudy (e.mccormick):

If I did this on graph paper it would help.... 

OpenStudy (e.mccormick):

Then I do the same thing for the other line.

OpenStudy (e.mccormick):

Can you find the intercepts for the other line? \( y=-\frac{3}{5}x+3\) ?

OpenStudy (anonymous):

No do help me out and also indicate the numbers on the graph

OpenStudy (e.mccormick):

Well, that is why you need some graph paper. Otherwise it is hard to tell where they cross.

OpenStudy (e.mccormick):

Whatever method you like for graphing lines is fine. I like intercepts a lot of the time. But if you graph some other way, it will work. just do one line, draw it in, then do the other, and draw it in, and see where they cross.

OpenStudy (anonymous):